eeeeeeeeeeeeeeee:

The equation below shows the total volume (V), in cubic units, of 5 identical boxes with each side equal to s units: V = 5s3 If s = 1.4 units, what is the value of V?

XioGonz:

Picture?

eeeeeeeeeeeeeeee:

what picture?

eeeeeeeeeeeeeeee:

13.72 cubic units 19.21 cubic units 21 cubic units 25 cubic units are the options

Convert:

V=5s3 V=5(1.4)(3) V=7(3) V=21

eeeeeeeeeeeeeeee:

ok! thanks!

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @eeeeeeeeeeeeeeee ok! thanks! \(\color{#0cbb34}{\text{End of Quote}}\) No problem! :)

Aricuzyas:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert V=5s3 V=5(1.4)(3) V=7(3) V=21 \(\color{#0cbb34}{\text{End of Quote}}\) Thanks @Convert